### v2

# Model-based Deep Learning for Joint RIS Phase Shift Compression and WMMSE Beamforming

Abstract
A model-based deep learning (DL) architecture is proposed for reconfigurable intelligent surface (RIS)-assisted multi-user communications to reduce the overhead of transmitting phase shift information from the access point (AP) to the RIS controller. The phase shifts are computed at the AP, which has access to the channel state information, and then encoded into a compressed binary control message that is sent to the RIS controller for element configuration. To help reduce beamformer mismatches due to phase shift compression errors, the beamformer is updated using weighted minimum mean square error (WMMSE) based on the effective channel resulting from the actual (decompressed) RIS reflection coefficients. By unrolling the iterative WMMSE algorithm as part of the wireless communication informed DL architecture, joint phase shift compression and WMMSE beamforming can be trained end-to-end. Simulations show that accounting for phase shift compression errors during beamforming significantly improves the sum-rate performance, even when the number of control bits is lower than the number of RIS elements.
I Introduction
The reconfigurable intelligent surface (RIS), a two-dimensional surface composed of several passive reflective elements, has attracted significant interest as a means to control the wireless propagation environment [1]. By jointly optimizing the RIS phase shifts and the beamformer, we can extend the signal coverage and enable cooperative signal focusing and interference cancellation along the wireless links between the access point (AP) and user equipments (UEs). This optimization takes place typically at the AP, which, having access to the channel state information (CSI), computes and transmits the phase shifts to the RIS controller.
Given its importance, joint optimization of the beamformers and RIS phase shifts is an active research topic. Various methods have been proposed to tackle this problem [2, 3, 4, 5, 6]. In particular, [2] proposed an alternating optimization (AO) method to minimize the transmit power, [3] proposed a block coordinate descent (BCD) method to maximize the weighted sum-rate, [4] proposed a weighted minimum mean square error power iteration (WMMSE-PI) method with lower complexity than the previous methods, [5] proposed a deep quantization neural network (DQNN) for the optimization of discrete phase shifts for each RIS element, and [6] proposed a deep learning (DL) architecture ACFNet that compresses and decompresses the bits of the RIS phase shifts between the AP and RIS controller. With the exception of [6], most studies overlook the phase shift control process, specifically, the overhead involved in transferring information from the AP to the RIS controller.
RISs are designed to be cost-effective and lack computing and sensing capabilities. As a result, computationally intensive tasks such as channel estimation [7, 8, 9, 10], and joint optimization of the beamforming and phase shifts are performed at the AP. The RIS controller, which updates the phase shifts, therefore relies entirely on the AP for phase shift reconfiguration. However, transmitting the phase shift value for each RIS element every time reconfiguration is required introduces a communication overhead that scales linearly with the number of elements, becoming a significant bottleneck in systems involving RISs with numerous elements [11]. This necessitates compression of the phase shifts which inevitably introduces decompression errors and a mismatch between the phase shifts assumed by the beamformer and those actually applied by the RIS which degrades the signal quality. Therefore, joint beamforming and phase shift optimization must account for the information transfer constraints between the AP and RIS controller.
Other than the ACFNet [6], previous works do not consider the phase shift control message transmission constraints. Despite the ACFNet compressing the RIS phase shifts, this DL model does not consider the effective channel with the decompressed phase shifts when jointly optimizing the beamforming and phase shifts. Since the ACFNet does not account for this mismatch, its compression gains come at the cost of significantly decreasing the achievable sum-rate. This performance gap between conventional iterative algorithms [2, 3, 4] that were designed by analyzing the wireless communication system model, versus previous generic DL methods that try to automatically discover the model information but result in overfitting [5, 6], indicates that there is room for improvement when designing a DL architecture with the phase shift control message transmission constraints.
In this paper, we propose the Auto-Quantization-Encoder and WMMSE beamforming updater (AQE-WMMSE) DL architecture to jointly optimize the AP beamformer and RIS phase shifts under a constraint on the number of control bits sent from the AP to the RIS controller. Specifically, our contributions are as follows:
  • We propose a joint beamforming and RIS phase shift DL architecture that compresses the phase shift information into a single control message. Compared to previous works [2, 3, 4, 5, 6], instead of individually transmitting phase shifts for each RIS element from the AP to the RIS controller, the proposed architecture takes advantage of the correlations between the phase shifts to extract and combine relevant features into the control message to reduce communication overhead.
  • Compared to conventional iterative algorithms [2, 3, 4], we jointly optimize the phase shifts and beamformer, maximizing the achievable sum-rate over multiple UEs, using end-to-end DL training. This is made possible by designing the AQE-WMMSE to use a trainable version of the iterative weighted minimum mean square error (WMMSE) algorithm based on unrolling principles [12].
  • The AQE-WMMSE accounts for the phase shift mismatch problem by explicitly integrating the compression process into the joint beamforming and phase shift optimization, and due to its model-based structure has a small hypothesis space which helps to avoid overfitting. In contrast, the generic neural networks (NNs) used in [5, 6] are prone to overfitting due to the generalized hypothesis space.
This paper is structured as follows. Section II describes the system model and the problem formulation, in Section III we propose a DL architecture and evaluate the proposed method against previous works in Section IV, finally we conclude the paper in Section V.
II System Model and Problem Statement
II-A System Model
We consider the downlink of a narrowband RIS-assisted communication system, as depicted in Fig. 1, comprising an AP with $M$ transmit antennas serving $K$ UEs each with one receive antenna, and an RIS with $N$ passive phase shift elements as depicted in Fig. 1.
Figure 1: RIS-assisted communication system.
The channels from the AP to the RIS, from the RIS to the $k$-th UE, and from the AP to the $k$-th UE are denoted by $\textbf{H}_{AR}\in\mathbb{C}^{N\times M}$, $\textbf{h}_{RU,k}\in\mathbb{C}^{1\times N}$, and $\textbf{h}_{AU,k}\in\mathbb{C}^{1\times M}$, respectively. All channels are assumed to be quasi-static and perfectly known [^1] Various techniques for channel estimation are explored in [7], [8], [9], and [10]. Assuming perfect CSI allows us to investigate upper bounds on the performance of RIS-assisted communications. at the AP. We also define the aggregate channel matrices $\textbf{H}_{RU}=[\textbf{h}_{RU,1}^{T},\ldots,\textbf{h}_{RU,K}^{T}]^{T}$, and $\textbf{H}_{AU}=[\textbf{h}_{AU,1}^{T},\ldots,\textbf{h}_{AU,K}^{T}]^{T}$. The AP beamforming vector for the $k$-th UE is denoted by $\textbf{w}_{k}\in\mathbb{C}^{M\times 1}$. Collectively, all $\textbf{w}_{k}$ satisfy a transmit power constraint $\sum_{k=1}^{K}\|\textbf{w}_{k}\|^{2}\leq P$, or $\|\textbf{W}\|^{2}_{F}\leq P$, where $\textbf{W}=[\textbf{w}_{1},\ldots,\textbf{w}_{K}]\in\mathbb{C}^{M\times K}$ is the beamforming matrix, and $\|\cdot\|_{F}$ denotes the Frobenius norm. The RIS reflection coefficients are represented by the vector $\boldsymbol{\phi}=[e^{j\theta_{1}},\ldots,e^{j\theta_{N}}]\in\mathbb{C}^{1\times N}$, where each $\theta_{n}\in[-\pi,\pi)$ denotes the phase shift applied by the $n$-th element, $1\leq n\leq N$. The phase shifts are collected in the vector $\boldsymbol{\theta}=[\theta_{1},\ldots,\theta_{N}]$.
The received signal $y_{k}\in\mathbb{C}$ at the $k$-th UE, due to the simultaneous transmission of the data symbols $s_{l}\in\mathbb{C}$, $1\leq l\leq K$, to the UEs is:
$y_{k}$ $=\sum_{l=1}^{K}(\textbf{h}_{AU,k}+{\boldsymbol{\phi}\textbf{H}_{k}})\textbf{w}_{l}s_{l}+n_{k}$ (1)
where $\textbf{H}_{k}=\operatorname{diag}({\textbf{h}_{RU,k}})\textbf{H}_{AR}\in\mathbb{C}^{N\times M}$ is the cascaded channel from the AP to the $k$-th UE via the RIS, and $n_{k}\sim\mathcal{CN}(0,\sigma^{2}_{k})$ is the additive white gaussian noise (AWGN) at the $k$-th UE. The achievable rate to the $k$-th UE is:
$\mathcal{R}_{k}=\text{log}_{2}\left(1+\frac{|(\textbf{h}_{AU,k}+\boldsymbol{\phi}\textbf{H}_{k})\textbf{w}_{k}|^{2}}{\sum_{l=1,l\neq k}^{K}|(\textbf{h}_{AU,l}+\boldsymbol{\phi}\textbf{H}_{l})\textbf{w}_{l}|^{2}+\sigma_{k}^{2}}\right)$ (2)
II-B Problem Statement
We assume that the AP has access to information $\mathcal{I}$ (such as the instantaneous CSI and the desired continuous RIS phase shifts), but the RIS controller does not have access to $\mathcal{I}$. To configure the RIS phase shifts, the AP computes and sends a binary control message, constrained to $B$ bits, to the RIS controller through a dedicated and error-free control link, separate from the main communication system. The controller decodes this message to determine the phase shifts used by the RIS. Formally, the process is defined by two functions: the AP generates the binary control message $\boldsymbol{\psi}\in\mathcal{F}=\{\textit{0},\textit{1}\}^{B}$ using $\boldsymbol{\psi}=f_{\text{c}}(\mathcal{I})$, while the RIS controller applies $\boldsymbol{\theta}=f_{\text{d}}(\boldsymbol{\psi})$ to decode $\boldsymbol{\psi}$ and obtains the phase shifts $\boldsymbol{\theta}$. In addition, the AP computes the beamforming matrix $\mathbf{W}$ from $\mathcal{I}$ using a third function $\mathbf{W}=f_{\text{w}}(\mathcal{I})$.
The problem we consider in this work is to identify $f_{\text{c}}(\cdot)$, $f_{\text{w}}(\cdot)$, and $f_{\text{d}}(\cdot)$ that jointly maximize the achievable sum-rate over the $K$ UEs. This problem can be formulated as follows:
$(\text{P1}):\quad\max_{f_{\text{c}}(\cdot),f_{\text{d}}(\cdot),f_{\text{w}}(\cdot)}\quad$ $\sum_{k=1}^{K}p_{k}\mathcal{R}_{k}$
s.t. $\boldsymbol{\psi}=f_{\text{c}}({\mathcal{I}})\in\mathcal{F}=\{\textit{0},\textit{1}\}^{B}$
$\boldsymbol{\theta}=f_{\text{d}}(\boldsymbol{\psi})$
$\textbf{W}=f_{\text{w}}({\mathcal{I}})\text{ with }\|\textbf{W}\|^{2}_{F}\leq P$ (3)
where $p_{k}$ is the priority of serving the $k$-th UE.
We note that, in the ideal case when the control link is not bandwidth-limited (i.e., $B\to\infty$), problem (P1) reduces to the traditional joint beamforming and RIS phase shift optimization problem. In this scenario, $\textbf{W}_{\text{opt}}=f_{\text{w}}({\mathcal{I}})$ and $\boldsymbol{\theta}_{\text{opt}}=f_{\text{d}}(f_{\text{c}}({\mathcal{I}}))$, where $(\textbf{W}_{\text{opt}},\boldsymbol{\theta}_{\text{opt}})$ denotes the optimal solution that can be obtained by AO [2], BCD [3], or WMMSE-PI [4].
III Proposed DL Architecture
In previous DL works [5, 6], generic NNs are used to solve (P1). In particular, a single NN combines the functions for beamforming $f_{\text{w}}(\cdot)$ and control message $f_{\text{c}}(\cdot)$ as $(\textbf{W},\boldsymbol{\psi})=f_{\text{w,c}}(\mathcal{I})$. Then for [6], a second NN $\boldsymbol{\theta}=f_{\text{d}}(\boldsymbol{\psi})$ decodes the control message $\boldsymbol{\psi}$. However, training these NNs $f_{\text{w,c}}(\cdot)$ and $f_{\text{d}}(\cdot)$ to obtain the solution $(\textbf{W},\boldsymbol{\theta})$ to (P1) can be challenging. This is because $f_{\text{w,c}}(\cdot)$, which solves for W, does not directly depend on the decoded phase shifts in $\boldsymbol{\theta}$, making the NN’s parameters have a large hypothesis space, resulting in overfitting [13, Section 5.2]. Instead, the beamforming matrix is updated indirectly with information about the phase shifts via backpropagation of a loss function $\mathcal{L}(\textbf{W},\boldsymbol{\theta})$ (see [5, eq. (5) and (13)] and [6, eq. (7) and (10)]). That is, the only way the beamforming matrix can be updated based on the phase shifts, is through the aggregate error of the batch of training samples and backpropagation of the corresponding gradients.
Figure 2: DL architecture of the AQE-WMMSE network.
In this paper, we propose to design the beamforming solution to explicitly depend on the phase shift values selected by the RIS controller. To avoid overfitting, we propose a wireless communication informed DL architecture AQE-WMMSE  [^2] Source code: https://github.com/AlexanderFernandes96/AQE-WMMSE to solve (P1) with two modules, as shown in Fig. 2. The first module: Auto-Quantization-Encoder (AQE), which is responsible for encoding the available information into a binary-valued control message, then decoding the control message into the RIS phase shifts. The second module: WMMSE beamforming updater, is responsible for updating the beamforming matrix with the actual RIS reflection coefficients by using the phase shifts at the output of the AQE. [^3] We note that an identical copy of the RIS controller’s AQE decoder is used at the AP to update W. Instead of using a generic NN as in [5, 6], to update the beamforming matrix, we propose a model-based approach by unrolling the WMMSE algorithm as part of our DL architecture. The input of the AQE-WMMSE is all the information available at the AP, $\mathcal{I}=\{\textbf{W}_{\text{opt}},\boldsymbol{\theta}_{\text{opt}},\textbf{H}_{AU},\textbf{H}_{AR},\textbf{H}_{RU}\}$, where ($\textbf{W}_{\text{opt}},\boldsymbol{\theta}_{\text{opt}})$ is the solution to (P1), obtained with the WMMSE-PI algorithm [4]. Both the WMMSE beamforming updater and the AQE modules are trained end-to-end using $\mathcal{I}$, which can be expressed in terms of solving (P1) as: $\boldsymbol{\theta}=f_{\text{d}}(f_{\text{c}}(\mathcal{I}))$ and $\textbf{W}=f_{\text{w}}(\mathcal{I};\boldsymbol{\theta})$. The notation $\textbf{W}=f_{\text{w}}(\cdot;\boldsymbol{\theta})$ signifies that W depends on $\boldsymbol{\theta}$.
III-A Auto-Quantization-Encoder
Similar to an autoencoder NN, the AQE network in Fig. 2 consists of an encoder part $f_{\text{EN}}(\cdot)$ and a decoder part $f_{\text{DE}}(\cdot)$, which are both implemented by NNs. In addition, the AQE uses an intermediate quantization layer $f_{\text{Q}}(\cdot)$ between the encoder and decoder NNs [14, 15]. In relation to (P1), the NNs perform the operations $\boldsymbol{\psi}=f_{\text{c}}(\mathcal{I})=f_{\text{Q}}(f_{\text{EN}}(\mathcal{I}))$ and $\boldsymbol{\theta}=f_{\text{d}}(\boldsymbol{\psi})=f_{\text{DE}}(\boldsymbol{\psi})$.
III-A1 Encoder
The NN $f_{\text{EN}}(\cdot)$ extracts $N_{c}$ phase features from the available information, $\tilde{\boldsymbol{\psi}}=[\tilde{\psi}_{1},\ldots,\tilde{\psi}_{Nc}]=f_{\text{EN}}(\mathcal{I})$.
III-A2 Quantizer
[…]
$\psi_{n}=Q(\tilde{\psi}_{n})=\sum_{i=1}^{D-1}a_{i}q(c_{i}(\tilde{\psi}_{n}-b_{i}))$ (4)
[…]
III-A3 Decoder
The decoder $f_{\text{DE}}(\cdot)$ reconstructs the $N$ RIS phase shifts in $\boldsymbol{\theta}$ from the control message $\boldsymbol{\psi}$.
III-B WMMSE Beamforming Updater
To solve (P1), we update $\textbf{W}_{\text{opt}}$ with $\boldsymbol{\theta}$ from the AQE, i.e. $\textbf{W}=f_{\text{w}}(\mathcal{I};\boldsymbol{\theta})$. To that end, we use the WMMSE algorithm [16] on the effective channel with the RIS phase shifts:
G $=\textbf{H}_{AU}+\textbf{H}_{RU}\operatorname{diag}(e^{j\boldsymbol{\theta}})\textbf{H}_{AR}$ (5)
where $\textbf{G}^{T}=[\textbf{g}_{1},\ldots,\textbf{g}_{K}]^{T}\in\mathbb{C}^{M\times K}$ for all $K$ UEs given $\boldsymbol{\theta}$. The WMMSE solution for the AP beamforming vector of the $k$-th UE from [16, eq. (15)] is:
$\textbf{w}_{k}$ $=p_{k}u_{k}\lambda_{k}\left(\sum_{l=1}^{K}p_{l}|u_{l}|^{2}\lambda_{l}\left(\frac{\sigma^{2}}{P}\textbf{I}_{M}+\textbf{g}_{l}^{H}\textbf{g}_{l}\right)\right)^{-1}\textbf{g}_{k}^{H}$ (6)
where $\textbf{u}=[u_{1},\ldots,u_{K}]\in\mathbb{C}^{K}$, and $\boldsymbol{\lambda}=[\lambda_{1},\ldots,\lambda_{K}]\in\mathbb{R}_{+}^{K},\lambda_{k}\geq 0$ are the WMMSE receiver and weight parameters, obtained iteratively using an alternating optimization approach given $\textbf{W}_{\text{opt}}$ as an initial value.
To train the AQE-WMMSE end-to-end, we unroll the iterative algorithm in [16] into the WMMSE beamforming updater depicted in Figure 2. We first propose the NN: $\textbf{w}^{\prime}=f_{w^{\prime}}(\textbf{W}_{\text{opt}},\textbf{G})$ as an initialization stage, where the vector $\textbf{w}^{\prime}\in\mathbb{R}^{2MK}$ represents updating the optimal beamforming matrix $\textbf{W}_{\text{opt}}$ based on the actual channel G. Then we propose the NN: $\textbf{u},\boldsymbol{\lambda}=f_{u\lambda}(\textbf{w}^{\prime},\textbf{G})$ to learn the WMMSE receiver and weight parameters [16, eq. (5) and (13)]. All together, $f_{w^{\prime}}$ initializes the beamformer, while $f_{u\lambda}$ and (6) implement a single iteration of the unrolled WMMSE algorithm (i.e. repeating $f_{u\lambda}$ and (6) as additional layers in the DL architecture could replicate multiple iterations of [16, Table 1]). Finally, the WMMSE beamforming matrix is $\textbf{W}=\frac{\sqrt{P}}{\sqrt{\sum_{k=1}^{K}\|\textbf{w}_{k}\|^{2}}}[\textbf{w}_{1},\ldots,\textbf{w}_{K}]\in\mathbb{C}^{M\times K}$ which satisfies the transmit power constraint.
IV Simulation Results
[…]
IV-A DL Parameters
The blocks [^4] Since current DL libraries such as PyTorch are unable to use complex values for NNs, all complex variables are represented as separate real and imaginary components, doubling the number of elements: $\textbf{u}\in\mathbb{C}^{K}\to\mathbb{R}^{2K}$. used in the proposed AQE-WMMSE architecture are structured as follows.
$f_{\text{EN}}$: Similar to [5], denoting $H=N+2MK$, we use a fully connected (FC) NN consisting of five linear layers of $32H$, $16H$, $8H$, $4H$, and $N_{c}$ neurons, respectively. The first four layers use the rectified linear unit (ReLU) activation function, while the last layer has no activation function. Between each pair of consecutive linear layers, we apply batch normalization followed by dropout with a probability of $0.5$.
$f_{\text{Q}}$: The scalar quantizer is set to 1-bit, i.e., $D=2$. Consequently, each encoder output $\psi_{n}$ corresponds to one bit of the control message, making the total number of control bits $B=N_{c}$. The amplitude $a_{1}$ is a non-trainable parameter set to $\frac{\pi}{2}$ resulting in two quantization levels: $\pm\frac{\pi}{2}$. The trainable parameters are initialized as $b_{1}=0$, and $c_{1}=0.5$.
$f_{\text{DE}}$: The decoder [^5] Since the decoder NN is implemented at the RIS controller, the decoder NN is designed to require less computation than the encoder NN at the AP. We also note that it is more difficult to update the parameters in the decoder than in the encoder, due to the vanishing gradients when passing information through the flat regions of the quantization layer [13]. is an FC NN with three linear layers each having $N$ neurons. The first two layers use ReLU, while the output layer has no activation function.
$f_{w^{\prime}}$ and $f_{u\lambda}$: both are four-layer FC NNs. Each of the first three layers has $4MK$ neurons (equal to the number of inputs) with a ReLU activation function followed by a dropout layer ($0.5$ probability). The last layer of $f_{w^{\prime}}$ and $f_{u\lambda}$ comprise $2MK$ and $3K$ neurons, respectively. No activation function is applied to all $2MK$ outputs of $f_{w^{\prime}}$ and $2K$ outputs of $f_{u\lambda}$, while the absolute value function is applied to the remaining $K$ outputs of $f_{u\lambda}$ to ensure non-negative elements for $\boldsymbol{\lambda}$.
IV-B Training Procedure
We generate a dataset of size $100,000$ and use $64,000$, $16,000$, and $20,000$ samples for training, validation, and testing, respectively. In all experiments, we use a batch size of $L=128$, and an early stopping strategy with patience of 50 to reduce overfitting, with the maximum number of epochs set to $1,000$. The learning rate is initialized to $0.001$ and is reduced by a factor of $0.8$ whenever the validation loss does not decrease for 20 consecutive epochs, until it reaches a minimum value of $0.00005$. The training loss function for all models is:
${\mathcal{L}}(\textbf{W},\boldsymbol{\theta})=-\frac{1}{S}\sum_{l=1}^{L}\sum_{k=1}^{K}p_{k}\mathcal{R}_{k}$ (7)
where $S$ is the sample size. For fair comparison to previous work, we use a priority $p_{k}=1$ for all $K$ UEs.
IV-C Benchmarks for Comparison
Upper Bound: we use the WMMSE-PI algorithm [4] with 100 iterations, to obtain the optimal AP beamforming matrix $\textbf{W}_{\text{opt}}$ and continuous valued RIS phase shifts $\boldsymbol{\theta}_{\text{opt}}$.
Benchmark 1 (DQNN [5]): a DL network that quantizes the phase shifts to $\pm\frac{\pi}{2}$, with each quantized value encoded using one bit. Therefore, the resulting control message has length $B=N$. The same DL network is also used to obtain the beamforming matrix based on the CSI [5, eq. (10), Fig. 2].
Benchmark 2 (ACFNet [6]): a DL network that compresses the bits of the control message. The DNN layer from [5, Fig. 2] is used as the input of the ACFNet method with $32H$, $16H$, $8H$, $4H$, and $2M+N$ neurons ($H=N+2MK$), in [6, Fig. 2] to obtain the beamforming matrix and phase shift vector before compression. For a fair comparison, compression is set to $B=N_{c}$ control bits by removing the Policy Network in [6, eq. (8) and Fig. 2] and the phase shifts at the RIS are continuous valued by removing $Q_{2}$ in [6, eq. (9) and Fig. 2].
Benchmark 3 (linQ [14]): a linear quantization method that updates the beamformer and compresses the bits of the control message. Given $\textbf{W}_{\text{opt}}$ and $\boldsymbol{\theta}_{\text{opt}}$, we implement linear analog combining and quantization [14] (i.e., NNs are matrix multiplication with bias and no activation function) to solve (P1), where $\boldsymbol{\theta}=f_{\text{D}}(f_{\text{Q}}(f_{\text{C}}(\boldsymbol{\theta}_{\text{opt}})))$ and $\textbf{W}=f_{W}(\textbf{W}_{\text{opt}})$. The linear operations ($f_{\text{C}}$, $f_{\text{D}}$, $f_{W}$) compress/decompress the phase shifts and update the beamformer.
All benchmarks use the same quantization scheme $f_{\text{Q}}(\cdot)$ outlined in Section IV-A to construct the control message. In the DQNN, ACFNet, and linQ benchmark methods, the beamforming matrix W is not calculated as a function of the actual phase shifts $\boldsymbol{\theta}$ chosen by the RIS controller.
IV-D Results
Figure 3: RIS phase shift control message fixed to $B=100$ bits (number of RIS elements) in (a) and compressed in (b) and (c).
In Fig. 3, we compare the achievable sum-rate for the proposed method with and without the WMMSE beamforming updater module (AQE-WMMSE and AQE) against the benchmarks. In Figure 3(a), we show the performance when the number of bits is fixed to be the number of RIS elements $B=N_{c}=N=100$. We also present the performance of each method when the control message fails to reach the RIS controller, in which case the RIS uses random phase shifts. Figures 3(b) and 3(c) compare the compression performance of each method where the number of bits of the control message becomes lower than the number of RIS elements, $B=N_{c}\leq N=100$. We can see that for high transmit power, AQE-WMMSE maintains a high achievable sum rate when compressing down from 100 bits to 40 bits and outperforms other methods even when compressing down to 10 bits.
(a) $P=15$ dBm
(b) $P=35$ dBm
Figure 4: Training/Validation loss for $B=40$ control bits.
Fig. 4 shows the AQE-WMMSE to have a lower validation loss (shorter curves are due to early stopping). This indicates that updating beamforming with the effective channel of the actual reflection coefficients can help improve the process in end-to-end training of the DL architecture.
V Conclusion
A wireless communication informed DL architecture was proposed to jointly compress the RIS phase shifts and update the AP beamforming matrix by reconstructing the phase shift information into a control message (transmitted between the AP and RIS controller). Numerical results showed that updating the beamforming matrix based on the effective channel (with the actual reflection coefficients of the RIS) obtained a higher achievable sum rate than recent literature methods when compressing the number of bits in the control message to be lower than the number of RIS elements.